OpenStudy (anonymous):

Let f(x) = x + 9 and g(x) = x – 3. What is f(x) • g(x)?

OpenStudy (anonymous):

x^2+6x-27

OpenStudy (anonymous):

(x+9)(x-3) =x^2-3x+9x-27 =x^2+6x-27!

OpenStudy (anonymous):

thanks!